Can you give the realation between the refrative index of a light of certain wavelength?? I mean formula or something.

Grade:Select Grade

1 Answers

Sumit Majumdar IIT Delhi
8 years ago
This is given by the standard Cauchy relation:
n(\lambda) = B + \frac {C}{\lambda^2} + \frac{D}{\lambda^4} + \cdots,
wherenis the refractive index, λ is the wavelength,B,C,D, etc., arecoefficientsthat can be determined for a material by fitting the equation to measured refractive indices at known wavelengths. The coefficients are usually quoted for λ as thevacuum wavelengthinmicrometres.

Usually, it is sufficient to use a two-term form of the equation:
n(\lambda) = B + \frac{C}{\lambda^2}
where the coefficientsBandCare determined specifically for this form of the equation.